OpenStudy (anonymous):

PLZ HLP MEDAL stuck on this one ... Use the given conditions to write an equation for the line in point-slope form. Passing through (-5, -7) and (-8, -6)

OpenStudy (ivettef365):

first find slope

OpenStudy (anonymous):

im reading up on textbooks n just not gettig this rightt now getting so aggervated with myself

OpenStudy (anonymous):

@mathstudent55

OpenStudy (ivettef365):

to find slope use formula y2 - y1 ------- x2 - x1 do you know how to ?

OpenStudy (anonymous):

no 6 years ago in high school i did lol

OpenStudy (ivettef365):

ok you have points (-5,-7) and (-8,-6) x1 y1 x2 y2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

tnx for ur help once again i apprciate it ! :)

OpenStudy (ivettef365):

that is just the first step, after you get the slope you need to use formula y = mx + b where m = slope and b = y-int

OpenStudy (anonymous):

-81/8 i got?.

OpenStudy (anonymous):

y + 6 = ^- 1/3(x + 8) ?

OpenStudy (ivettef365):

that is correct

OpenStudy (anonymous):

yay! thank you